the following items deal with play ground swing: a) how could you increase its frequency of oscillation? b) at which point(s) in the swing are you accelerating linearly and why? c) where along your swing path is your speed the greatest? d) would this swing work in deep space ,explain?
Answer #1

T = 2*π*sqrt(L/g)
f = 1/T
f = 1/(2*π) * sqrt(g/L)

To increase Freq, Length of the swing need to be deceased.

(b)
A swing has positive linear acceleration when it falls.

(c)
Using Energy Conservation,
At the bottom of the path, P.E is minimum and K.E is maximum,
Therefore Spped of swing is Greatest at bottom of the path.

(d)
No it won't as Gravitational Acceleration (g) is negligible at deep space, and it is because of gravity the swing works.